OpenStudy (anonymous):

if f(x)=3x^2+9x-10 find and simplify f(2+x)

OpenStudy (anonymous):

Put x=2+x in place of x in given equation & solve it

OpenStudy (anonymous):

so 3(2+x)^2+9(2+x)-10?

OpenStudy (anonymous):

Yes Just proceed with this

OpenStudy (anonymous):

do i square everything the the parentesees?

OpenStudy (anonymous):

ok i got 3x^2+21x-20

OpenStudy (anonymous):

Use (x+y)^2= x^2+y^2+2*x*y

OpenStudy (anonymous):

i got 3x^2+21x+20 after i simplified is that correct?

OpenStudy (anonymous):

It will be +20 instead of -20

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

Welcome :)